The sum of three numbers is one hundred fifteen. The second number is three times the first number. The third number is two less than five times the first number. Find the three numbers
Found 2 solutions by xcentaur, Homeworkboy_06:
Answer by xcentaur(357)

let the numbers be x,y,z


then x+y+z=115 (given)
also y=3x (given)
also z=(5x-2) (given)


x+y+z=115
x+3x+(5x-2)=115
9x=115+2
x=117/9=13


x=13
y=3x=39
z=5x-2=65-2=63


13,39,63


Prabhat


Answer by Homeworkboy_06(1)